George can rake a lawn in 5 hours. George and Juan together can rake the lawn in 2 hours. How long would it take for Juan to rake the lawn by himself? Use the following equation to solve this problem where J represents the time it would take Juan to rake the lawn:
2/5 + 2/J =1

Answer by nerdybill(7384)

George can rake a lawn in 5 hours. George and Juan together can rake the lawn in 2 hours. How long would it take for Juan to rake the lawn by himself? Use the following equation to solve this problem where J represents the time it would take Juan to rake the lawn:
2/5 + 2/J =1
multiplying both sides by 5J:
2J + 10 = 5J
10 = 3J
10/3 hours = J
or
3 hours and 20 minutes = J